OpenStudy (anonymous):

Question: -Label and display your new polynomial identity -Prove that it is true through an algebraic proof, identifying each step -Demonstrate that your polynomial identity works on numerical relationships Below is a list of some sample factors you may use to help develop your own identity. (x – y) (x + y) (y + x) (y – x) (x + a) (y + b) (x2 + 2xy + y2) (x2 – 2xy + y2) (ax + b) (cy + d) *** I do not need a straight answer, I need guidance. For example, I do not know which is the best set of factors to chose.

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

haven't learned polynomial yet

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

so it sounds like you have to use those factors and multiply them out you don't have to choose all of them, just 2 or 3 maybe

jimthompson5910 (jim_thompson5910):

(x – y)*(x+y) = ???

OpenStudy (anonymous):

I'm not allowed to use that order because it was used in the lesson

OpenStudy (anonymous):

Would (y+x)(y-x) work?

jimthompson5910 (jim_thompson5910):

sure that works

jimthompson5910 (jim_thompson5910):

what does that multiply out to?

OpenStudy (anonymous):

FOIL right?

OpenStudy (anonymous):

y^2-xy+xy-x^2?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

can you simplify that at all?

OpenStudy (anonymous):

Yep! do the -xy and +xy cancel out or do they become -2xy? If so it would be y^2-2xy-x^2

jimthompson5910 (jim_thompson5910):

nope, more like 0xy since the -1 and the +1 add to 0

jimthompson5910 (jim_thompson5910):

it would be -2xy if you had -xy-xy

OpenStudy (anonymous):

Thats what I was thinking

jimthompson5910 (jim_thompson5910):

0xy is the same as 0, so it goes away

OpenStudy (anonymous):

Okay then so it would be y^2-x^2?

jimthompson5910 (jim_thompson5910):

yes it is

jimthompson5910 (jim_thompson5910):

(y+x)(y-x) = y^2 - x^2 is an identity

OpenStudy (anonymous):

Awesome!

jimthompson5910 (jim_thompson5910):

an identity is any equation that is always true for the allowed input variables

OpenStudy (anonymous):

What's next, how would I prove it

jimthompson5910 (jim_thompson5910):

well you effectively did really when you used the distributive property, that is an algebraic property

jimthompson5910 (jim_thompson5910):

when you FOIL, you're really using the distributive property twice

OpenStudy (anonymous):

Is there a name to this identity?

OpenStudy (anonymous):

Okay I see

jimthompson5910 (jim_thompson5910):

(y+x)(y-x) = y^2 - x^2 difference of squares

OpenStudy (anonymous):

Ohhh okay cool! So when the question states: Demonstrate that your polynomial identity works on numerical relationships. What does that mean?

jimthompson5910 (jim_thompson5910):

you just plug in various x,y pairs

jimthompson5910 (jim_thompson5910):

and prove the equation holds true

OpenStudy (anonymous):

And the numbers that I plug in can be random?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

And I would plug in those numbers to the identity that we just made? to both sides?

jimthompson5910 (jim_thompson5910):

For example (x,y) = (7,2) ------------- (y+x)(y-x) = y^2 - x^2 (2+7)(2-7) = 2^2 - 7^2 (9)(-5) = 4 - 49 -45 = -45

jimthompson5910 (jim_thompson5910):

so (y+x)(y-x) = y^2 - x^2 has been proven true for the specific case when x = 7, y = 2

OpenStudy (anonymous):

Wow great example! I understand now

jimthompson5910 (jim_thompson5910):

what's another example?

OpenStudy (anonymous):

Okay here is another one: (x,y) = (2,4) (y+x)(y-x) = y^2 - x^2 (4+2)(4-2) = 4^2 - 2^2 (6)(2)=16-4 12=12

OpenStudy (anonymous):

Is that correct or did I make a mistake?

jimthompson5910 (jim_thompson5910):

you nailed it, nice work

jimthompson5910 (jim_thompson5910):

so if you managed to do EVERY possible x,y pair allowed, then you'd find that they all work. Of course, there are infinitely many pairs, so it's not possible to do this. This is why the algebraic proof is much more effective

OpenStudy (anonymous):

Yay!!! So the whole thing is done?

OpenStudy (anonymous):

I really appreciate you helping me

jimthompson5910 (jim_thompson5910):

yes it is and I'm glad I could help